A 100g ball and 200g ball are connected by a rod with a length of 60 cm. The mass of the rod is ignored. The axis of rotation is located at the center of the rod. What is the moment of inertia of the balls about axis of rotation?


1
Expert's answer
2021-12-03T12:44:48-0500
"I=(m+M)(0.5l)^2\\I=(0.1+0.2)(0.3)^2=0.027\ kgm^2"